# MnLargeSymbols’164 MnLargeSymbols’171 MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
An appropriate choice of batch sizes in large-scale model training is crucial, yet it involves an intrinsic yet inevitable dilemma: large-batch training improves training efficiency in terms of memory utilization, while generalization performance often deteriorates due to small amounts of gradient noise. Despite this dilemma, the common practice of choosing batch sizes in language model training often prioritizes training efficiency—employing either constant large sizes with data parallelism or implementing batch size warmup schedules. However, such batch size schedule designs remain heuristic and often fail to adapt to training dynamics, presenting the challenge of designing adaptive batch size schedules. Given the abundance of available datasets and the data-hungry nature of language models, data parallelism has become an indispensable distributed training paradigm, enabling the use of larger batch sizes for gradient computation. However, vanilla data parallelism requires replicas of model parameters, gradients, and optimizer states at each worker, which prohibits training larger models with billions of parameters. To optimize memory usage, more advanced parallelism strategies must be employed. In this work, we propose general-purpose and theoretically principled adaptive batch size schedules compatible with data parallelism and model parallelism. We develop a practical implementation with PyTorch Fully Sharded Data Parallel, facilitating the pretraining of language models of different sizes. We empirically demonstrate that our proposed approaches outperform constant batch sizes and heuristic batch size warmup schedules in the pretraining of models in the Llama family, with particular focus on smaller models with up to 3 billion parameters. We also establish theoretical convergence guarantees for such adaptive batch size schedules with Adam for general smooth nonconvex objectives.
@@ 1 Introduction
Large-batch training (i.e., using large batch sizes) is arguably the current de facto training paradigm for large language models, driven by recent advances and the availability of computational hardware for deep learning. For instance, the open-sourced model, OLMo 65B (Groeneveld et al., 2024), utilizes a batch size of 2048 sequences of length 2048, resulting in 4M tokens per batch. Despite the efficient utilization of available hardware through parallelization, a major drawback of large-batch training is the issue of “generalization gap” (see e.g., LeCun et al., 2002)—where model generalization performance deteriorates compared to small-batch training without heavy tuning of other hyperparameters. See Figure 1 for a graphical illustration of the existence of generalization gaps with different batch sizes when training a vanilla transformer with 61M parameters. Keskar et al. (2017) argued that small-batch methods tend to converge to flat minima, leading to better generalization. To close this generalization gap, several works (Hoffer et al., 2017; Smith et al., 2018; Smith and Le, 2018) have proposed using large learning rates to offset the effect of large batch sizes, recovering the generalization performance of using small batches. However, the training of language (and vision) models based on the attention mechanism (Vaswani et al., 2017) and the transformer architecture is notoriously unstable. Reducing training instability, including unwanted loss spikes (see e.g., Zhai et al., 2023; Wortsman et al., 2024), demands significant tuning and cautious hyperparameter selections, like using a small learning rate.
@@
Beyond using a large learning rate to balance the intrinsic trade-off between training efficiency and generalization performance of large-batch training, Keskar et al. (2017) also suggested the use of adaptive sampling methods (Byrd et al., 2012; Friedlander and Schmidt, 2012). These methods are essentially adaptive batch size schemes that progressively improve the accuracy of the batch gradient approximation by gradually increasing batch sizes throughout the model training process. This concept has been explored by De et al. (2016, 2017); Lau et al. (2024b), but their implementations are limited to the single-device setting, where all data samples are implicitly assumed to reside on the same device. This limitation makes them unfit for data-parallel distributed training wherein data is spread across various workers in a parallel system, potentially encompassing several network-connected nodes, thereby preventing the scaling necessary to train large models. Beyond the single-device setting, Lau et al. (2024a) have also extended such adaptive batch size schemes to local gradient methods for local batch sizes, where model synchronization is performed every several gradient steps rather than every step.
@@
Data parallelism (Krizhevsky et al., 2012), such as DistributedDataParallel (DDP) in PyTorch (Li et al., 2020) and counterparts in TensorFlow (Abadi et al., 2015) and JAX (Bradbury et al., 2018; Frostig et al., 2018), is arguably the most popular paradigm for distributed training in deep learning. In data parallelism (alone), each worker holds a local copy of the model parameters (as well as gradient and optimizer states). The global input batch is divided into multiple minibatches for each training step, so each worker performs forward and backward computations with a different minibatch. After each training step, all GPUs perform an all-reduce collective communication to synchronize gradients, followed by a global model parameter update. This ensures that all local copies of the model remain identical after the parameter update steps. Adaptive batch size schemes can be developed based on the approaches in Byrd et al. (2012); Friedlander and Schmidt (2012); Bollapragada et al. (2018) for data-parallel settings, providing practical adaptive batch size schedules in PyTorch DDP for training large-scale deep neural networks, which require data parallelism.
@@
While these practical schemes open up the possibility of distributed training of larger models with GPUs of lower memory, they are constrained by the inherent design of DDP—the need to maintain a model replica at each worker. State-of-the-art large language models (LLMs) now consist of billions or even hundreds of billions of parameters (e.g., Llama 3 405B; Llama Team, AI @ Meta, 2024). Distributed training with only data parallelism thus unfortunately fails, as the memory required to store such models well exceeds the available memory of a single GPU. Even worse, access to expensive workstation-level GPUs with more memory is often limited to industrial labs, whereas academic researchers and end-users often have to resort to less powerful consumer-level GPUs or workstation-level GPUs with less memory.
@@
To alleviate this limitation inherent to data parallelism, more memory-efficient paradigms of parallelism, such as model parallelism (Shoeybi et al., 2019), have been proposed. In model parallelism, model parameters are sharded into various components and distributed to different workers. In particular, PyTorch Fully Sharded Data Parallel (FSDP; Zhao et al., 2023) is an implementation of model parallelism in PyTorch (Paszke et al., 2019), marking the first native feature in PyTorch that can support models with up to trillions of parameters without relying on more sophisticated third-party libraries for model parallelism such as DeepSpeed (Rasley et al., 2020), Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021; Korthikanti et al., 2023), and their combinations (Smith et al., 2022), which could be overwhelming to get started with and too technical to modify for users’ specific needs. Moreover, PyTorch FSDP has been widely adopted in the pretraining of various open-source LLMs such as OPT (Zhang et al., 2022), TinyLlama (Zhang et al., 2024b), OLMo (Groeneveld et al., 2024), and DRBX (The Mosaic Research Team, 2024).
@@
However, even with data parallelism and model parallelism, LLM pretraining involving models with up to hundreds of billions of parameters and trillions of tokens (e.g., Llama 3 405B; Llama Team, AI @ Meta, 2024), still incurs extensive costs (more than millions of US dollars per model) and imposes a significant carbon footprint. Consequently, there is a pressing need for developing proper and well-crafted training strategies. In this work, we focus on choosing dynamic batch size schedules, which deserve more attention than they have, since, unlike other optimizer hyperparameters, batch sizes also control training efficiency via memory utilization of GPUs, in addition to affecting model generalization performance and training stability. The current practice of choosing batch sizes in LLM pretraining, however, remains heuristic, in the sense that it usually involves either constant large batch sizes or prespecified heuristic warmup schedules which could be very hard to design.
@@
In this work, we propose theoretically principled adaptive batch size schedules based on the adaptive sampling method (Byrd et al., 2012) for pretraining large language models, which are also generally applicable to training other deep neural networks. On the theoretical front, we establish a convergence guarantee for the proposed adaptive batch size schedules for Adam, the de facto optimizer for pretraining language models. Various recent works have shown, both empirically and theoretically, that Adam outperforms SGD in training attention-based language models (Pan and Li, 2023; Kunstner et al., 2023, 2024; Zhang et al., 2024c). Our convergence guarantee complements the existing results of adaptive batch size schedules for SGD (De et al., 2016, 2017) and AdaGrad (Lau et al., 2024b). From a practical perspective, we develop a solution of adaptive batch size schedules based on PyTorch FSDP, which are tailor-made for pretraining LLMs with more than billions of parameters.
@@ 2 Related Work
Large-batch training has proven to be very successful for different deep learning applications including computer vision (Goyal et al., 2017; Akiba et al., 2017) and natural language processing (You et al., 2020; Liu et al., 2019; Puri et al., 2018). From an empirical perspective, many open-source or open-weights models, such as OPT (Zhang et al., 2022), BLOOM (BigScience Workshop et al., 2022), Mistral 7B (Jiang et al., 2023), Baichuan 2 (Yang et al., 2023), Qwen (Bai et al., 2023; Yang et al., 2024), OLMo (Groeneveld et al., 2024), Gemma (Gemma Team, 2024; Gemma Team et al., 2024), Llama (Touvron et al., 2023; Llama Team, AI @ Meta, 2024) and DeepSeek (DeepSeek-AI et al., 2024a, b), revealed that they were pretrained with large numbers of GPUs or TPUs (i.e., data-parallel sizes), hence naturally making use of large-batch training. While using large batch sizes is now standard, the rationale for choosing the magnitude of such large batch sizes is mostly based on hardware availability. Only recently in the training of Stable LM 2 1.6B, Bellagente et al. (2024) clarified the selection of global batch sizes, aiming to strike an optimal balance between minimizing training time and the extra training tokens needed to reach the desired final training loss. Shallue et al. (2019) study the effects of data parallelism by performing ablation studies on different batch sizes by training different models on different datasets using different optimizers, finding no evidence that large batch sizes degrade generalization performance with careful hyperparameter search. From a more theoretical perspective, McCandlish et al. (2018) develop a model for understanding the critical batch size that determines the tradeoff between speed and efficiency of large-batch training. Kaplan et al. (2020) further study the scaling law of the critical batch size as a power of the training loss only for language models. However, in most of these works, benchmarking was performed with different magnitudes of constant batch sizes, with the notable exception of McCandlish et al. (2018) which provided a case study of dynamically varying the batch size with an adaptive batch size schedule, but only using a simple model (CNN) and dataset (SVHN). The effect of adaptive batch sizes for pretraining language models, to the best of our knowledge, remains elusive to the community.
@@
Adaptive sampling methods (Byrd et al., 2012; Friedlander and Schmidt, 2012; Bollapragada et al., 2018), which adjust batch sizes based on gradient noise or gradient approximation quality, are further explored in deep learning (De et al., 2016, 2017; Lau et al., 2024b; Ostroukhov et al., 2024) but have not been applied to data parallelism with distributed samplers. The development of adaptive batch size schedules for deep learning is not a novel concept, featuring methodologies such as Big Batch SGD (De et al., 2016, 2017), CABS (Balles et al., 2017), AdaBatch (Devarakonda et al., 2017), SimiGrad (Qin et al., 2021) and AdaScale SGD (Johnson et al., 2020). Our work is also closely related to and motivated by the heuristic technique of batch size warmup/batch ramp, which has been widely adopted in pretraining LLMs and even in reinforcement learning (Hilton et al., 2023). Batch size warmup usually involves prespecified schedules of multiple batch size stages starting from training with multiple increasing smaller batch sizes for small portions of the total training tokens, followed by training with the remaining tokens using a large batch size. For instance, GPT-3 (Brown et al., 2020) was pretrained by gradually increasing the batch size linearly from a small value (32k tokens) to the full value (3.2M tokens) over the first 4–12 billion tokens of training. Nemotron-4 (Parmar et al., 2024) was pretrained with a batch size schedule of batch sizes 384–768–1152 sequences for 2.5%–2.5%–95% of the total number of training tokens. Likewise, OLMo 65B (Groeneveld et al., 2024) was pretrained with a batch size warmup schedule of batch sizes 2M–4M–8M–16M tokens, using 100B tokens for the first three stages and the largest batch size for the remaining training tokens (2T total training tokens). Llama 3 405B (Llama Team, AI @ Meta, 2024) was trained using the following batch size schedule: an initial batch size of 4M tokens with a sequence length 4096 tokens for 252M tokens; a batch size of 8M tokens with a sequence length of 8192 tokens for 2.87T tokens; a batch size of 16M tokens for the remainder of a total of about 15T training tokens. Such a batch size recipe is found to be able to stabilize training—few loss spikes were observed and it did not require interventions to correct for model training divergence. Despite potentially improving training efficiency or data parallelism, batch size warmup schedules remain heuristic and their impact on training is difficult to grasp. Another related yet seemingly orthogonal technique is sequence length warmup (Li et al., 2022; Jin et al., 2023), which progressively grows the sequence length throughout the pretraining process. Note that the pretraining of Llama 3 405B employs both batch size warmup and sequence length warmup.
@@ 3.1 Vanilla Adaptive Batch Size Schedules
where $\ell\colon\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}\cup\{\pm\infty\}$ is the individual loss function, and $\mathcal{D}_{n}\coloneqq\{z_{i}\}_{i=1}^{n}$ is the set of $n$ training samples. If $\ell(\cdot;z)$ is continuously differentiable for any $z\in\mathcal{Z}$, then the gradient of the loss function and its batch counterpart (i.e., the batch gradient) are given by $\nabla\mathscr{L}(w)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(w;z_{i})$ and $\nabla\mathscr{L}_{\mathcal{B}}(w)\coloneqq\frac{1}{b}\sum_{i\in\mathcal{B}}\nabla\ell(w;z_{i})$, where the batch $\mathcal{B}\subseteq\llbracket n\rrbracket$ is a subset of indices of data points sampled uniformly without replacement, and $b\coloneqq|\mathcal{B}|$ is the corresponding batch size. We write $\ell_{i}(w)\coloneqq\ell(w;z_{i})$ and $\nabla\ell_{i}(w)\coloneqq\nabla\ell(w;z_{i})$. The batch gradient $\nabla\mathscr{L}_{\mathcal{B}}$ is used to approximate the full gradient $\nabla\mathscr{L}$ as the number of samples $n$ is prohibitively large.
@@
Falling into the family of adaptive sampling methods, the norm test (Byrd et al., 2012) is motivated by measuring the quality of the approximation of the full gradient $\nabla\mathscr{L}$ by the batch gradient $\nabla\mathscr{L}_{\mathcal{B}}$ through the lens of approximation of a descent direction for the loss $\mathscr{L}$. If $\ell(\cdot;z)$ is also convex, then $-\nabla\mathscr{L}_{\mathcal{B}}$ is a descent direction for $\mathscr{L}$ at $w\in\mathbb{R}^{d}$ if and only if $\langle\mathscr{L}_{\mathcal{B}}(w),\mathscr{L}(w)\rangle\geqslant 0$, that is, $\mathscr{L}_{\mathcal{B}}$ and $\mathscr{L}$ share the same direction at $w$. It can be shown that the above inner product condition is equivalent to the norm condition: $\left\lVert\nabla\mathscr{L}_{\mathcal{B}}(w)-\nabla\mathscr{L}(w)\right\rVert\leqslant\eta\left\lVert\nabla\mathscr{L}(w)\right\rVert$ for any $\eta\in[0,1)$. This condition cannot be checked directly, since the number of samples $n$ is in billions for LLMs and the full gradient $\nabla\mathscr{L}$ is unavailable. Instead, we have to resort to a batch approximation:
@@
where $\mathrm{Var}_{i\in\mathcal{B}}(\nabla\ell_{i}(w))\coloneqq\frac{1}{b-1}\sum_{i\in\mathcal{B}}\left(\nabla\ell_{i}(w)-\nabla\mathscr{L}_{\mathcal{B}}(w)\right)^{2}$. The adjustment factor $(n-b)/(n-1)$ is approximated by $1$ as we take $n\to\infty$. Consequently, to ensure that the batch gradient approximates the descent direction of the full objective $\mathscr{L}$ well, the (approximate) norm test checks the following condition at each iteration $k\in\mathbb{N}^{*}$:

### v2

@@ Abstract
An appropriate choice of batch sizes in large-scale model training is crucial, yet it involves an intrinsic yet inevitable dilemma: large-batch training improves training efficiency in terms of memory utilization, while generalization performance often deteriorates due to small amounts of gradient noise. Despite this dilemma, the common practice of choosing batch sizes in language model training often prioritizes training efficiency—employing either constant large sizes with data parallelism or implementing batch size warmup schedules. However, such batch size schedule designs remain heuristic and often fail to adapt to training dynamics, presenting the challenge of designing adaptive batch size schedules. Given the abundance of available datasets and the data-hungry nature of language models, data parallelism has become an indispensable distributed training paradigm, enabling the use of larger batch sizes for gradient computation. However, vanilla data parallelism requires replicas of model parameters, gradients, and optimizer states at each worker, which prohibits training larger models with billions of parameters. To optimize memory usage, more advanced parallelism strategies must be employed. In this work, we propose general-purpose and theoretically principled adaptive batch size schedules compatible with data parallelism and model parallelism. We develop a practical implementation with PyTorch Fully Sharded Data Parallel, facilitating the pretraining of language models of different sizes. We empirically demonstrate that our proposed approaches outperform constant batch sizes and heuristic batch size warmup schedules in the pretraining of models in the Llama 2 family, with particular focus on smaller models with up to 3 billion parameters. We also establish theoretical convergence guarantees for such adaptive batch size schedules with Adam for general smooth nonconvex objectives.
@@ 1 Introduction
Large-batch training (i.e., using large batch sizes) is arguably the current de facto training paradigm for large language models, driven by recent advances and the availability of computational hardware for deep learning. For instance, the open-weight model, Llama 3 405B [50], utilizes a batch size of 1024 sequences of length 4096, resulting in 4M tokens per batch. Despite the efficient utilization of available hardware through parallelization, a major drawback of large-batch training is the issue of “generalization gap” (see e.g., [43])—where model generalization performance deteriorates compared to small-batch training without heavy tuning of other hyperparameters. See Figure 1 for a graphical illustration of the existence of generalization gaps with different batch sizes when training a vanilla transformer with 61M parameters. Keskar et al. [34] argued that small-batch methods tend to converge to flat minima, leading to better generalization. To close this generalization gap, several works [29, 68, 67] have proposed using large learning rates to offset the effect of large batch sizes, recovering the generalization performance of using small batches. However, the training of language (and vision) models based on the attention mechanism [74] and the transformer architecture is notoriously unstable. Reducing training instability, including unwanted loss spikes (see e.g., [82, 78]), demands significant tuning and cautious hyperparameter selections, like using a small learning rate.
@@
Beyond using a large learning rate to balance the intrinsic trade-off between training efficiency and generalization performance of large-batch training, Keskar et al. [34] also suggested the use of adaptive sampling methods [12, 21]. These methods are essentially adaptive batch size schemes that progressively improve the accuracy of the batch gradient approximation by gradually increasing batch sizes throughout the model training process. This concept has been explored by De et al. [13, 14] and Lau et al. [42], but their implementations are limited to the single-device setting, where all data samples are implicitly assumed to reside on the same device. This limitation makes them unfit for data-parallel distributed training wherein data is spread across various workers in a parallel system, potentially encompassing several network-connected nodes, thereby preventing the scaling necessary to train large models. Beyond the single-device setting, Lau et al. [41] have also extended such adaptive batch size schemes to local gradient methods for local batch sizes, where model synchronization is performed every several gradient steps rather than every step.
@@
Data parallelism [38], such as DistributedDataParallel (DDP) in PyTorch [46] and counterparts in TensorFlow [1] and JAX [10, 22], is arguably the most popular paradigm for distributed training in deep learning. In data parallelism (alone), each worker holds a local copy of the model parameters (as well as gradient and optimizer states). The global input batch is divided into multiple minibatches for each training step, so each worker performs forward and backward computations with a different minibatch. After each training step, all GPUs perform an all-reduce collective communication to synchronize gradients, followed by a global model parameter update. This ensures that all local copies of the model remain identical after the parameter update steps. Adaptive batch size schemes can be developed based on the approaches in [12, 21, 8] for data-parallel settings, providing practical adaptive batch size schedules in PyTorch DDP for training large-scale deep neural networks, which require data parallelism.
@@
While these practical schemes open up the possibility of distributed training of larger models with GPUs of lower memory, they are constrained by the inherent design of DDP—the need to maintain a model replica at each worker. State-of-the-art large language models (LLMs) now consist of billions or even hundreds of billions of parameters (e.g., Llama 3 405B [50]). Distributed training with only data parallelism thus unfortunately fails, as the memory required to store such models well exceeds the available memory of a single GPU. Even worse, access to expensive workstation-level GPUs with more memory is often limited to industrial labs, whereas academic researchers and end-users often have to resort to less powerful consumer-level GPUs or workstation-level GPUs with less memory.
@@
To alleviate this limitation inherent to data parallelism, more memory-efficient paradigms of parallelism, such as model parallelism [66], have been proposed. In model parallelism, model parameters are sharded into various components and distributed to different workers. In particular, PyTorch Fully Sharded Data Parallel (FSDP) [88] is an implementation of model parallelism in PyTorch [57], marking the first native feature in PyTorch that can support models with up to trillions of parameters without relying on more sophisticated third-party libraries for model parallelism such as DeepSpeed [62], Megatron-LM [66, 53, 37], and their combinations [69], which could be overwhelming to get started with and too technical to modify for users’ specific needs. Moreover, PyTorch FSDP has been widely adopted in the pretraining of various open-source language models such as OPT [85], TinyLlama [84], OLMo [27, 71], and DRBX [72].
@@
However, even with data parallelism and model parallelism, LLM pretraining involving models with up to hundreds of billions of parameters and trillions of tokens (e.g., Llama 3 405B [50]), still incurs extensive costs (more than millions of US dollars per model) and imposes a significant carbon footprint. Consequently, there is a pressing need for developing proper and well-crafted training strategies. In this work, we focus on choosing dynamic batch size schedules, which deserve more attention than they have, since, unlike other optimizer hyperparameters, batch sizes also control training efficiency via memory utilization of GPUs, in addition to affecting model generalization performance and training stability. The current practice of choosing batch sizes in LLM pretraining, however, remains heuristic, in the sense that it usually involves either constant large batch sizes or prespecified heuristic warmup schedules which could be very hard to design.
@@
In this work, we propose theoretically principled adaptive batch size schedules based on the adaptive sampling method [12] for pretraining large language models, which are also generally applicable to training other deep neural networks. On the theoretical front, we establish a convergence guarantee for the proposed adaptive batch size schedules for Adam, the de facto optimizer for pretraining language models. Various recent works have shown, both empirically and theoretically, that Adam outperforms SGD in training attention-based language models [55, 39, 40, 86]. Our convergence guarantee complements the existing results of adaptive batch size schedules for SGD [13, 14] and AdaGrad [42]. From a practical perspective, we develop a solution of adaptive batch size schedules based on PyTorch FSDP, which are tailor-made for pretraining LLMs with more than billions of parameters.
@@ 2 Related Work
Large-batch training has proven to be very successful for different deep learning applications including computer vision [26, 2] and natural language processing [81, 49, 58]. From an empirical perspective, many open-source or open-weights models, such as OPT [85], BLOOM [6], Mistral 7B [30], Baichuan 2 [79], Qwen [3, 80], OLMo [27, 71], Gemma [24, 23], Llama [73, 50] and DeepSeek [15, 16], revealed that they were pretrained with large numbers of GPUs or TPUs (i.e., data-parallel sizes), hence naturally making use of large-batch training. While using large batch sizes is now standard, the rationale for choosing the magnitude of such large batch sizes is mostly based on hardware availability. Only recently in the training of Stable LM 2 1.6B, Bellagente et al. [5] clarified the selection of global batch sizes, aiming to strike an optimal balance between minimizing training time and the extra training tokens needed to reach the desired final training loss. Shallue et al. [65] study the effects of data parallelism by performing ablation studies on different batch sizes by training different models on different datasets using different optimizers, finding no evidence that large batch sizes degrade generalization performance with careful hyperparameter search. From a more theoretical perspective, McCandlish et al. [52] develop a model for understanding the critical batch size that determines the tradeoff between speed and efficiency of large-batch training. Kaplan et al. [33] further study the scaling law of the critical batch size as a power of the training loss only for language models. However, in most of these works, benchmarking was performed with different magnitudes of constant batch sizes, with the notable exception of McCandlish et al. [52] which provided a case study of dynamically varying the batch size with an adaptive batch size schedule, but only using a simple model (CNN) and dataset (SVHN). The effect of adaptive batch sizes for pretraining language models, to the best of our knowledge, remains elusive to the community.
@@
Adaptive sampling methods [12, 21, 8], which adjust batch sizes based on gradient noise or gradient approximation quality, are further explored in deep learning [13, 14, 42, 54] but have not been applied to data parallelism with distributed samplers. The development of adaptive batch size schedules for deep learning is not a novel concept, featuring methodologies such as Big Batch SGD [13, 14], CABS [4], AdaBatch [17], SimiGrad [59] and AdaScale SGD [32]. Our work is also closely related to and motivated by the heuristic technique of batch size warmup/batch ramp, which has been widely adopted in pretraining LLMs and even in reinforcement learning [28]. Batch size warmup usually involves prespecified schedules of multiple batch size stages starting from training with multiple increasing smaller batch sizes for small portions of the total training tokens, followed by training with the remaining tokens using a large batch size. For instance, GPT-3 [11] was pretrained by gradually increasing the batch size linearly from a small value (32k tokens) to the full value (3.2M tokens) over the first 4–12 billion tokens of training. Nemotron-4 [56] was pretrained with a batch size schedule of batch sizes 384–768–1152 sequences for 2.5%–2.5%–95% of the total number of training tokens. Llama 3 405B [50] was trained using the following batch size schedule: an initial batch size of 4M tokens with a sequence length 4096 tokens for 252M tokens; a batch size of 8M tokens with a sequence length of 8192 tokens for 2.87T tokens; a batch size of 16M tokens for the remainder of a total of about 15T training tokens. Such a batch size recipe is found to be able to stabilize training—few loss spikes were observed and it did not require interventions to correct for model training divergence. Despite potentially improving training efficiency or data parallelism, batch size warmup schedules remain heuristic and their impact on training is difficult to grasp. Another related yet seemingly orthogonal technique is sequence length warmup [45, 31], which progressively grows the sequence length throughout the pretraining process. Note that the pretraining of Llama 3 405B employs both batch size warmup and sequence length warmup.
@@ 3.1 Vanilla Adaptive Batch Size Schedules
where $\ell\colon\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}\cup\{\pm\infty\}$ is the individual loss function, and $\mathcal{D}_{n}\coloneqq\{z_{i}\}_{i=1}^{n}$ is the set of $n$ training samples. If $\ell(\cdot;z)$ is continuously differentiable for any $z\in\mathcal{Z}$, then the gradient of the loss function and its batch counterpart (i.e., the batch gradient) are given by
$\nabla\mathscr{L}(w)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(w;z_{i})\quad\text{and}\quad\nabla\mathscr{L}_{\mathcal{B}}(w)\coloneqq\frac{1}{b}\sum_{i\in\mathcal{B}}\nabla\ell(w;z_{i}),$
where the batch $\mathcal{B}\subseteq\llbracket n\rrbracket$ is a subset of indices of data points sampled uniformly without replacement, and $b\coloneqq|\mathcal{B}|$ is the corresponding batch size. We write $\ell_{i}(w)\coloneqq\ell(w;z_{i})$ and $\nabla\ell_{i}(w)\coloneqq\nabla\ell(w;z_{i})$. The batch gradient $\nabla\mathscr{L}_{\mathcal{B}}$ is used to approximate the full gradient $\nabla\mathscr{L}$ as the number of samples $n$ is prohibitively large.
@@
Falling into the family of adaptive sampling methods, the norm test [12] is motivated by measuring the quality of the approximation of the full gradient $\nabla\mathscr{L}$ by the batch gradient $\nabla\mathscr{L}_{\mathcal{B}}$ through the lens of approximation of a descent direction for the loss $\mathscr{L}$. If $\ell(\cdot;z)$ is also convex, then $-\nabla\mathscr{L}_{\mathcal{B}}$ is a descent direction for $\mathscr{L}$ at $w\in\mathbb{R}^{d}$ if and only if $\langle\mathscr{L}_{\mathcal{B}}(w),\mathscr{L}(w)\rangle\geqslant 0$, that is, $\mathscr{L}_{\mathcal{B}}$ and $\mathscr{L}$ share the same direction at $w$. It can be shown that the above inner product condition is equivalent to the norm condition: $\left\lVert\nabla\mathscr{L}_{\mathcal{B}}(w)-\nabla\mathscr{L}(w)\right\rVert\leqslant\eta\left\lVert\nabla\mathscr{L}(w)\right\rVert$ for any $\eta\in[0,1)$. This condition cannot be checked directly, since the number of samples $n$ is in billions for LLMs and the full gradient $\nabla\mathscr{L}$ is unavailable. Instead, we have to resort to a batch approximation:
@@
where
$\mathrm{Var}_{i\in\mathcal{B}}(\nabla\ell_{i}(w))\coloneqq\frac{1}{b-1}\sum_{i\in\mathcal{B}}\left(\nabla\ell_{i}(w)-\nabla\mathscr{L}_{\mathcal{B}}(w)\right)^{2}.$
The adjustment factor $(n-b)/(n-1)$ is approximated by $1$ as we take $n\to\infty$. Consequently, to ensure that the batch gradient approximates the descent direction of the full objective $\mathscr{L}$ well, the (approximate) norm test checks the following condition at each iteration $k\in\mathbb{N}^{*}$: